South Pole Diaries 1997/98

   

   


Sunday 30th November 1997 - Alice's Restaurant

From John Storey....

On the day after Thanksgiving the cooks take the day off, so this year CARA and JACARA stepped into the breech to provide brunch for the station. This consisted mainly of omlettes expertly cooked by Jim, while the rest of us ran around cutting up leftovers and opening tins of things that go in omlettes. Vegemite on crackers was provided for those feeling a bit delicate after the previous night, and the whole event rocked along to a CD of The Sixties Down-under: the quintessential sampling of the best rock music a young nation could deliver 30 years ago. The brunch patrons seemed equally divided on the what was worse - the Vegemite or the music.

Brunch concluded with a full volume performance of Alice's Restaurant - to the delight of all those over 40 and the bemused wonderment of all those under.

Dinner was provided by the "ASTRO" team. "Ten-metre Tony's" beef-in-beer stew (option for vegetarians: leave out the beef) was, frankly, a bit ordinary, but serves well to prepare me for the horrors of the McMurdo canteen next week. Ant and Al did a great job of washing dishes etc. - well beyond the call of duty.

Last night I attempted to run the MISM for the first time since we'd reassembled it. Nothing would work and I was unable to telnet in. A quick investigation showed that one of the inductors in the power supply was getting very hot - which is surprising because, as every physicist knows, an ideal inductor cannot dissipate energy. Either this inductor had decided to choose this moment to become seriously non-ideal, or it was being forced to deal with a much higher current than it liked. The latter theory was soon confirmed by the observation that the 5V supply was being dragged down to 3.5V, which was why nothing was working. I considered this to be such an unsatisfactory state of affairs that I went to bed.

Today we took another look at it, and it quickly became clear that the 5V line out to the optics box was shorted to ground. Now, who has been fiddling with the optics box recently?, I hear you ask. Sure enough, we'd succeeded in bolting the lid down in such a way that it crushed the wires - a suprisingly easy thing to do once your fingers are frozen stiff and given that you can't see the stupid wires anyway. Nothing for it but to put the optics plate off again, warm it up inside in a plastic bag, and make yet another repair. This time we made the wires *really* short - for two reasons: one, so they wouldn't get caught in the lid again, and two, that was all that was left of them.

The fact that the Maxim power supply was able to merrily thump 3 amps into the short, while at the same time raising the magnetic flux in the inductor to dizzy heights way beyond its wildest dreams, confirms what a fabulous piece of technology it is. Tip for today: buy shares in Maxim.

Anyway, with everything back together again, the Mism *appears* to be 100% up and running. Ant even stood on a ladder and waved a soldering iron around to confirm that both beams were there.

There's still a couple of idiosyncracies of the software that make it difficult to go much further. For example, "choptest" reports the correct frequency (nearly - I've no idea how it can be off by 5% when the same computer is generating *and* measuring the frequency) for the *first* sample only. Then, either the chopper decides to spontaneously drop its frequency by 10Hz, or the software is off with the faries:

choptest Jan 01 03:18:35 CHOPPER_TEST Jan 01 03:18:35 allowable percentage error? > .1 .1 Jan 01 03:18:37 total length of test? (seconds) > 10 10 making 5 iterations of 2000 samples; each iteration takes 2 seconds, each sample takes 1.00 milliseconds hit a key to abort 1, ref = 1000.000, actual = 995.243, delta = -4.757, stdev = 2020.704 2, ref = 1000.000, actual = 986.054, delta = -13.946, stdev = 2054.730 3, ref = 1000.000, actual = 986.385, delta = -13.615, stdev = 2056.661 4, ref = 1000.000, actual = 985.540, delta = -14.460, stdev = 2057.502 5, ref = 1000.000, actual = 985.613, delta = -14.387, stdev = 2055.884

Next, attempting to use "sample" generates the following output: sample 100 Jan 01 03:24:19 TAKE_DATA_SAMPLES value = 0.000; stdev = 0.000; count = 2; channel = 0 Error #2195 - STS_HWF_PHASELOCK_2 (3)

I take this to mean that the computer is not seeing the 2xfrequency interrupts. But does "choptest" use these interrupts, or the 1xfrequency square wave? I'll check tomorrow to make sure that all the correct waveforms are actually going into the computer. Note that all the above is with the chopper well and truly phase locked.

The vaccuum in Abu is great (3 E-6 in whatever units these things are measured in - problably pounds per square inch), and Al and Ant are preparing to fire up the closed-cycle cooler and do the first South-pole cooldown. They've got as far as determining that they need to change a plug or two, or call in an electrician. Unfortunately, electricians are a bit thin on the ground at the moment as two of them were booted off the station last weekend after they had a punch-up. This is a misdemeanour that results in automatic expulsion. By all acoounts it was just a bloke thing. For reasons I've never understood, some blokes, after they'd had a skinfull, instead of becoming warm and fuzzy and at peace with the world prefer to belt the tripe out of each other as a mark of affection. Something to do with testosterone, I think.

Al Fowler claims you can freeze cockroaches in liquid nitrogen, and when they thaw out they just get up and walk around. He says they're good for at least three cooldown cycles before they start to get a bit wobbly. Can anyone corroborate this? If not, it sounds like a good honours project - with maybe expanding the range of experimental animals up to and including cats (and possibly first-year students).

Michael A. has been running the nism, and found that the chopper was very slow to phase-lock. This is a bit odd, as it was doing fine the other day. When I tried it today, sure enough it was very marginal. A small tweak on the centre-frequency pot brought it up to par, but why it should have changed since the other day I do not know. I'll keep an eye on it. I also had trouble today getting the "start nism" command to work, but as soon as my friend the Tektronics Digital Sampling CRO stepped in to take a look, the problem instantly vanished.

Today was a little bit more windy, but the station is still totally peaceful. With no aircraft, no bulldozers, and almost no people outside, for the first time one gets to feel how isolated it is here. A few people came out to jog around the skiway (in preparation for the "Round-the World race), while others took a stroll along the skiway - an activity best reserved for days on which there are no flights. Occasionally a beaker would go scudding past in the Sprite, but apart from the muffled sound of the diesel power generator all was still.

Tomorrow the station will spring back to life again. We're hoping Dave Pernic will arrive - he's one of those thoroughly useful people who can do anything, including make the mounts for the SPIREX secondary mirror.

Our plan is to have Abu on the telescope by Friday. Everyone says that this is completely unrealistic, but no-one is sure why.

John

 

 

Information

Further Information

Contact: